OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)? y = mx + b m is the slope b is the y-intercept

OpenStudy (anonymous):

@ameenbriscoe13 @Firefly007 @fallenangelorchid @humphriescody @jojo4eva @KamiBug @purplepiggy171 @Xd1521 @vickerys37 @horsegirl27

OpenStudy (horsegirl27):

i can help just let me finish something else

OpenStudy (anonymous):

@horsegirl27 ok

OpenStudy (anonymous):

Sorry math isn't really my thing.

OpenStudy (anonymous):

im so sorry i suck at math i could figure it out but i dont see any picture :( sorry

OpenStudy (anonymous):

@vickerys37 @Firefly007 that's ok

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

sory I do not now

OpenStudy (anonymous):

@humphriescody ok :(

OpenStudy (fallenangelorchid):

Multiple choice...

OpenStudy (anonymous):

no you have to find the answer for the question no multiple choice sorry

OpenStudy (fallenangelorchid):

kk

OpenStudy (horsegirl27):

first find the slope

OpenStudy (anonymous):

ive tried the equation multiple times I really don't know what the slope is

OpenStudy (horsegirl27):

i think 2

OpenStudy (horsegirl27):

is there any more to this question? I'm not sure of the y intercept normally it's given to you

OpenStudy (anonymous):

no that's the whole question

myininaya (myininaya):

slope can be found by doing \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{14-6}{9-5}\]

OpenStudy (horsegirl27):

oh I know what it is

myininaya (myininaya):

\[y=mx+b \] after you find the slope we can use one the points to find b

OpenStudy (horsegirl27):

@myininaya thats what I did. Did I do it right getting 2

OpenStudy (anonymous):

wow thanks @myininaya

OpenStudy (horsegirl27):

is the y intercept 14? I feel like I did this before and got 14...

myininaya (myininaya):

yes 8/4=2 so m=2 so you have y=mx+b where m is 2 so we actually have so far y=2x+b you find b by using either point plug in (5,6) to find b that solve the following equaiton for b: 6=2(5)+b

OpenStudy (anonymous):

ok thanks so much I was so confused @myininaya

OpenStudy (anonymous):

so b is -4 :D

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

6-10=b yes b=-4

myininaya (myininaya):

so you have that you want the form y=mx+b and you were to just find m and b and you done that so you have m=2 and b=-4 your equation in slope intercept form is y=2x-4

myininaya (myininaya):

good job